OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9234 — Optical loop Lyon
Incident Report for Web Cloud
Resolved
We have a fibre down this morning in the Lyon optical loop, securisation has taken over. No outage expected. A technician is en route to carry out optic measurements.

Update(s):

Date: 2013-09-02 18:03:55 UTC
There was a broken fibre in an optic case. Everything is operational again, securisation is in place.

Date: 2013-09-02 16:10:58 UTC
The cut in the fibre of the operator has been detected 2500m from the SPR site, a team has been deployed to takae care of the soldering.

Date: 2013-09-02 12:26:25 UTC
We have carried out the measures on our side. We are no longer getting a signal on an optic strand between SPR and MPL. A ticket has been opened with our fibre provider.
Posted Sep 02, 2013 - 09:21 UTC